# STRICT: Stress Test of Rendering Images Containing Text

### v1

@@ Abstract
While diffusion models have revolutionized text-to-image generation with their ability to synthesize realistic and diverse scenes, they continue to struggle to generate consistent and legible text within images. This shortcoming is commonly attributed to the locality bias inherent in diffusion-based generation, which limits their ability to model long-range spatial dependencies. In this paper, we introduce STRICT, a benchmark designed to systematically stress-test the ability of diffusion models to render coherent and instruction-aligned text in images. Our benchmark evaluates models across multiple dimensions: (1) the maximum length of readable text that can be generated; (2) the correctness and legibility of the generated text, and (3) the ratio of not following instructions for generating text. We evaluat several state-of-the-art models, including proprietary and open-source variants, and reveal persistent limitations in long-range consistency and instruction-following capabilities. Our findings provide insights into architectural bottlenecks and motivate future research directions in multimodal generative modeling. We release our entire evaluation pipeline at https://github.com/tianyu-z/STRICT-Bench.
STRICT: Stress Test of Rendering Images Containing Text
Tianyu Zhang1,2*, Xinyu Wang3*, Zhenghan Tai4, Lu Li5, Jijun Chi4, Jingrui Tian6, Hailin He7, Suyuchen Wang1,2 1DIRO, Université de Montréal 2Mila - Quebec AI Institute 3McGill University 4University of Toronto 5University of Pennsylvania 6University of California, Los Angeles 7Southwestern University of Finance and Economics {tianyu.zhang, suyuchen.wang}@umontreal.ca, xinyu.wang5@mail.mcgill.ca, luli1@sas.upenn.edu, {winfred.tai, ferris.chi}@mail.utoronto.ca, nakotian@g.ucla.edu
@@ 1 Introduction
Recent advances, such as OpenAI’s Image-4o OpenAI (2025), have shown promising progress in this domain, achieving near-human performance in rendering embedded text. Similarly, open-source models such as HiDream-L1 HiDream.ai (2025) and SeedDream 3 Gao et al. (2025) have reported comparable success in overcoming long-range dependency issues. However, a systematic and quantitative evaluation of these capabilities remains lacking.
In this work, we present STRICT (Stress Test of Rendering Images Containing Text), a comprehensive benchmark designed to rigorously evaluate the performance of diffusion models in generating image-embedded text. Our contributions are threefold:
We introduce STRICT, a multilingual benchmark that tests model performance on rendering texts of varying lengths in English, Chinese, and French.
We propose quantitative metrics for assessing (1) the maximum readable text length, (2) the correctness of the generated content, and (3) the ratio of not following instructions for generating text in images.
@@
Through this evaluation, our objective is to expose current limitations, identify failure patterns, and guide the development of structure-aware generative models capable of producing semantically and visually coherent text within images.
@@ 2 Task Design
The primary objective of our benchmark is to rigorously evaluate the capability of text-to-image generation models to render accurate and coherent text embedded within synthetic images. The task is to utilize the model to be evaluated with a ground-truth content string and instructed with a natural language prompt template: “Produce an image of a typed document page with the following text: [TEXT]” to generate an image containing this specific text [TEXT]. The characteristics of the ground-truth text sample are systematically varied to probe different aspects of model performance. These variations include:
Text Length: Ground-truth texts range from short phrases to longer paragraphs to determine the maximum length of text a model can generate coherently. This helps assess how accuracy degrades as the quantity of text increases.
@@
For each sample, a model generates an image based on the input prompt containing the target ground-truth text, which then forms a pair with its corresponding ground-truth text contents. We collect these generated images containing embedded text to evaluate each model.
@@ 3 Evaluation Metrics
To quantify the performance of models in rendering text, we employ an OCR-based verification framework. The text content from each generated image is first extracted using an OCR engine, followed by a comparison against the original ground-truth text.
@@ 3.1 OCR and Preprocessing
We utilize the Tesseract OCR engine Ooms (2025) to extract text from generated images. Tesseract is a widely adopted open-source OCR tool capable of recognizing over 100 languages and offering configurable page segmentation modes (PSM) to suit different layout scenarios. In our evaluation, we primarily use English, French, and Chinese language models with PSM set to 3, which indicates fully automatic page layout analysis.
To ensure a fair and robust evaluation, we perform minimal but critical preprocessing of both the OCR-extracted text and the ground-truth references. Specifically, we adopt a strict text processing strategy that normalizes whitespace across both texts, collapsing all forms of whitespace (spaces, tabs, newlines) into a single space, and trims leading and trailing whitespace.
After preprocessing, we evaluate OCR performance using a suite of metrics: character-level accuracy (Character Error Rate, CER) (Radford et al., 2023; Conneau et al., 2021), word-level accuracy (Word Error Rate, WER) (Morris et al., 2004; Kim et al., 2021), and normalized edit distance (NED) (Fisman et al., 2022). For a more granular analysis, we compute these metrics under two settings: a full comparison, which compares the entire OCR output to the full ground truth, and a truncated comparison, which compares the OCR output against a truncated version of the ground truth matched to the number of tokens recognized by OCR. This dual-mode evaluation allows us to assess not only absolute performance but also the model’s ability to preserve textual order and correctness under realistic generation length constraints.
Through these evaluations, we reveal systematic differences in how text-to-image models perform under varying linguistic and spatial constraints in multilingual settings.
@@ 3.2 Quantitative Metrics
Character Error Rate (CER): Other than NED, we also use two commonly used metrics in speech recognition: Character Error Rate (CER) (Radford et al., 2023; Conneau et al., 2021) and Word Error Rate (WER) (Morris et al., 2004; Kim et al., 2021). Similar to NED, CER operates at the character level. It is more sensitive to minor errors like single incorrect letters or OCR misrecognitions. CER is also normalized by the number of characters in the ground truth (for full CER), with lower values being better. Note that the CER metric does not have an upper limit.
@@
Word Error Rate (WER): WER is a standard word-level metric commonly seen in speech recognition and OCR, which calculates the minimum number of word-level insertions, deletions, and substitutions required to transform the OCR output into ground truth text. The result is typically normalized by the number of words in the ground truth text (for Full WER). Lower WER values signify higher accuracy. Note that the WER metric does not have an upper limit.
@@ 4 Experiment Results
We evaluate a diverse set of state-of-the-art text-to-image generation models, including proprietary and open-source variants, on the STRICT benchmark. The models tested include: GPT-4o (gpt-image-1) OpenAI (2025), Seedream 3.0 Gao et al. (2025), Recraft V3 AI (2024a), HiDream-I1-Dev HiDream.ai (2025), Imagen 3 (imagen-3.0-generate-002) DeepMind (2024a), FLUX 1.1 pro Labs (2024), Gemini 2.0 (gemini-2.0-flash-preview-image-generation) DeepMind (2024b), as well as open-source models such as Stable Diffusion 3.5 Medium AI (2024b), Anytext 2 Zhao and Lian (2024a), and the TextDiffuser 2 Chen et al. (2023b, 2024). The text dataset is sourced from Wikipedia Foundation , which is licensed under the Creative Commons Attribution-ShareAlike 3.0 (CC-BY-SA 3.0) license. The OCR processing was performed using Tesseract OCR Ooms (2025), which is distributed under the Apache License 2.0. Both licenses permit use for research and commercial purposes, provided the respective attribution and license terms are followed.
@@
According to Figure 2, across all evaluated languages and text lengths, GPT-4o and Gemini-2.0 outperform all competing models significantly by a large margin in terms of character accuracy, word accuracy, and instruction adherence. We evaluated ten models with varying capabilities. For the weakest models (Anytext 2, TextDiffuser 2, Stable Diffusion 3.5 Medium), we tested input lengths ranging from 5 to 300 characters. For moderately performing models (FLUX 1.1 pro, Seedream 3.0, HiDream-I1-Dev), the tested character lengths ranged from 50 to 2,000. Although moderate performance, the API upper limit of Recraft V3 is 1000 bytes (1,000 Latin characters or 500 Chinese characters). Thus, we test Recraft V3 from 50 to its upper limits. Moreover, the strongest models (Imagen 3, Gemini 2.0 and GPT-4o) were evaluated on inputs ranging from 50 to 5,000 characters. For both English and French, GPT-4o and Gemini maintain strong performance up to approximately 800 characters, beyond which accuracy begins to decrease. Detailed CER and WER metrics are shown in Appendix B for reference. For Chinese, the overall model performance remains poor. However, GPT-4o still consistently outperforms the other models. For detailed scores, please check Table 1.
@@ 5 Discussion
Most diffusion models demonstrate a marked decline in performance as the input text length exceeds approximately 200 characters. This threshold aligns with the 77-token limit of the CLIP text encoder, which is used in HiDream, stable diffusion series and other diffusion models. Taking a step back, as mentioned in Zhang et al. (2024a), the real effective token length of CLIP is 20 rather than 77. Considering the above two points together, it is likely that the model’s ability to capture and condition on longer instructions is restricted. Moreover, existing training corpora rarely include such long and instruction-heavy prompts, further exacerbating generalization challenges in these situations.
@@
With longer text prompts, diffusion models increasingly fail to adhere to instruction semantics. Instead of generating a document-like image containing the target text, many models instead synthesize a naturalistic image related to the subject matter of the prompt, completely omitting any embedded text. For example, in Figure 4, when instructed to render a document stating “Asia ( , ) is the largest continent in the world by both land area and population…”, some models, especially Flux 1.1 pro, return an illustration of an Asian map instead of the text itself as the number of characters increase. We hypothesize that this failure arises from three main causes: (1) the dilution of the instruction signal in long prompts, which reduces the model’s focus on the instruction of “document should contain the text”; (2) the limited capacity of CLIP-based or similar text encoders to capture long-range logical structure in extended inputs. These encoders may process such inputs more like a bag-of-words and thus compromise instruction fidelity; (3) the use of certain models, such as HiDream-L1, which incorporate LLMs (e.g., Llama-3.1-8B Grattafiori et al. (2024)) as encoders, although LLMs are primarily designed as decoders and are not well-suited for encoding tasks. Previous works Huang et al. (2024); BehnamGhader et al. (2024) have highlighted that LLMs generally underperform in encoder roles, although some modifications can improve their effectiveness.
@@
We observe consistent differences in performance across languages. English generally yields the highest accuracy, followed by French, while Chinese exhibits the lowest performance. This is likely attributable to insufficient training data for Chinese, rather than intrinsic difficulties in rendering Chinese characters. Structurally, we do not think that there is a fundamental reason why showing Chinese characters should be harder to generate than generating the Latin alphabet. The performance gap underscores the need for broader multilingual data inclusion in models’ pretraining.
@@ Appendix A Appendix: Prompt Variants
Appendix B Appendix: CER and WER Figure

### v2

# STRICT: Stress-Test of Rendering Image Containing Text

@@ Abstract
While diffusion models have revolutionized text-to-image generation with their ability to synthesize realistic and diverse scenes, they continue to struggle with generating consistent and legible text within images. This shortcoming is commonly attributed to the locality bias inherent in diffusion-based generation, which limits their capacity to model long-range spatial dependencies. In this paper, we introduce STRICT, a benchmark designed to systematically stress-test the ability of diffusion models to render coherent and instruction-aligned text in images. Our benchmark evaluates models across multiple dimensions: (1) the maximum length of readable text that can be generated; (2) the correctness and legibility of the generated text, and (3) the ratio of not following instructions for generating text. We assess several state-of-the-art models, including proprietary and open-source variants, and reveal persistent limitations in long-range consistency and instruction-following capabilities. Our findings provide insights into architectural bottlenecks and motivate future research directions in multimodal generative modeling. We release all our evaluation pipeline at https://github.com/tianyu-z/STRICT-Bench/.
STRICT: Stress-Test of Rendering Image Containing Text
Tianyu Zhang1††thanks: These authors contributed equally, Xinyu Wang211footnotemark: 1, Lu Li311footnotemark: 1, Zhenghan Tai4, Jijun Chi4, Jingrui Tian5, Hailin He6 Suyuchen Wang1 1Mila, University of Montreal 2McGill University 3University of Pennsylvania 4University of Toronto 5University of California, Los Angeles 6Southwestern University of Finance and Economics
@@ 1 Introduction
Recent advances, such as OpenAI’s Image-4o OpenAI (2025), have shown promising progress in this domain, achieving near-human performance in rendering embedded text. Similarly, open-source models like HiDream-L1 HiDream.ai (2025) and SeedDream 3 Gao et al. (2025) have reported comparable success in overcoming long-range dependency issues. Yet, a systematic and quantitative evaluation of these capabilities remains lacking.
In this work, we present STRICT (Stress-Test of Rendering Image Containing Text), a comprehensive benchmark designed to rigorously evaluate the performance of diffusion models in generating image-embedded text. Our contributions are threefold:
We introduce a multi-lingual benchmark that tests model performance on rendering texts of varying lengths in English, Chinese, and French.
We propose quantitative metrics for assessing (1) the maximum readable text length; (2) the correctness of the generated content, and (3) the ratio of not following instructions for generating text.
@@
Through this evaluation, we aim to expose current limitations, identify failure patterns, and guide the development of structure-aware generative models capable of producing semantically and visually coherent text within images.
@@ 2 Task Design
The primary objective of our benchmark is to rigorously evaluate the capability of text-to-image generation models to render accurate and coherent text embedded within synthetic images. The task is to utilize the model to be evaluated with a ground truth content string and instructed with a natural language prompt template: f"Produce an image of a typed document page with the following text: [TEXT] " to generate an image containing this specific text [TEXT]. The characteristics of the ground truth text sample are systematically varied to probe different aspects of model performance. These variations include:
Text Length: Ground truth texts range from short phrases to longer paragraphs to determine the maximum length of text a model can generate coherently. This helps assess how accuracy degrades as the quantity of text increases.
@@
For each sample, a model generates an image based on the input prompt containing the target ground truth text and then forms a pair with its corresponding ground truth text file. We collect these generated image-text pairs for evaluating each model.
@@ 3 Evaluation Metrics
To quantify the performance of models in rendering text, we employ an OCR-based verification framework. The text content from each generated image is first extracted using an OCR engine, followed by a comparison against the original ground truth text.
@@ 3.1 OCR and Preprocessing
We utilize the Tesseract OCR engine Ooms (2025) for extracting text from generated images. Tesseract is a widely adopted open-source OCR tool capable of recognizing over 100 languages and offering configurable page segmentation modes (PSM) to suit different layout scenarios. In our evaluation, we primarily use English, French, and Chinese language models with PSM set to 3, which indicates fully automatic page layout analysis.
To ensure fair and robust evaluation, we perform minimal but critical preprocessing of both the OCR-extracted text and the ground truth references. Specifically, we adopt a strict text processing strategy that normalizes whitespace across both texts, collapsing all forms of whitespace (spaces, tabs, newlines) into a single space, and trims leading and trailing whitespace.
After preprocessing, we evaluate OCR performance using a suite of metrics: character-level accuracy (Character Error Rate, CER) (Radford et al., 2023; Conneau et al., 2021), word-level accuracy (Word Error Rate, WER) (Morris et al., 2004; Kim et al., 2021), normalized edit distance (NED) Fisman et al. (2022), and sequence similarity using Python package difflib Python Software Foundation (2001). For a more granular analysis, we compute these metrics under two settings: a full comparison, which compares the entire OCR output to the full ground truth, and a truncated comparison, which compares the OCR output against a truncated version of the ground truth matched to the number of tokens recognized by OCR. This dual-mode evaluation allows us to assess not only absolute performance but also the model’s ability to preserve textual order and correctness under realistic generation length constraints.
Through these evaluations, we reveal systematic differences in how text-to-image models perform under varying linguistic and spatial constraints in multi-language settings.
@@ 3.2 Quantitative Metrics
Character Error Rate (CER): Other than NED, we also use two commonly used metrics in speech recognition: Character Error Rate (CER) (Radford et al., 2023; Conneau et al., 2021) and Word Error Rate (WER) (Morris et al., 2004; Kim et al., 2021). Similar to NED, CER operates at the character level. It is more sensitive to minor errors like single incorrect letters or OCR misrecognitions. CER is also normalized by the number of characters in the ground truth (for Full CER), with lower values being better. Note that the CER values do not have an upper limit.
@@
Word Error Rate (WER): WER is a standard word-level metric commonly seen in speech recognition and OCR, calculating the minimum number of word-level insertions, deletions, and substitutions required to transform the OCR output into the ground truth text. The result is typically normalized by the number of words in the ground truth text (for Full WER). Lower WER values signify higher accuracy. Note that the WER values do not have an upper limit.
@@ 4 Experiment Results
We evaluate a diverse set of state-of-the-art text-to-image generation models, including proprietary and open-source variants, on the STRICT benchmark. The models tested include: GPT-4o (gpt-image-1) OpenAI (2025), Seedream 3.0 Gao et al. (2025), Recraft V3 AI (2024a), HiDream-I1-Dev HiDream.ai (2025), Imagen 3 (imagen-3.0-generate-002) DeepMind (2024a), FLUX 1.1 pro Labs (2024), Gemini 2.0 (gemini-2.0-flash-preview-image-generation) DeepMind (2024b), as well as open-source models such as Stable Diffusion 3.5 Medium AI (2024b), Anytext 2 Zhao and Lian (2024a), and the TextDiffuser 2 Chen et al. (2023b, 2024). We also include partial results for Qwen-Image Wu et al. (2025) and gemini-2.5-flash (nano-banana) DeepMind (2025), which is released after we submit the paper. The text dataset is sourced from Wikipedia Foundation , which is licensed under the Creative Commons Attribution-ShareAlike 3.0 (CC-BY-SA 3.0) license. The OCR processing was performed using Tesseract OCR Ooms (2025), which is distributed under the Apache License 2.0. Both licenses permit use for research and commercial purposes, provided the respective attribution and license terms are followed.
@@
According to Figure 3, across all evaluated languages and text length, GPT-4o and Gemini-2.0 outperform all competing models significantly by a large margin in terms of character accuracy, word accuracy, and instruction adherence. We evaluated ten models with varying capabilities. For the weakest models (Anytext 2, TextDiffuser 2, Stable Diffusion 3.5 Medium), we tested input lengths ranging from 5 to 300 characters. For moderately performing models (FLUX 1.1 pro, Seedream 3.0, HiDream-I1-Dev), the tested character lengths ranged from 50 to 2,000. Although moderate performance, the API upper limit of Recraft V3 is 1000 bytes (1,000 Latin characters or 500 Chinese characters). Thus, we test Recraft V3 from 50 to its upper limits. Moreover, the strongest models (Imagen 3, Gemini 2.0 and GPT-4o) were evaluated on inputs ranging from 50 to 5,000 characters. For both English and French, GPT-4o and Gemini maintain strong performance up to approximately 800 characters, beyond which accuracy begins to decrease. Detailed CER and WER metrics are displayed in the Appendix C for reference. For Chinese, the overall model performance remains poor. However, GPT-4o still consistently outperforms the other models. We display all the scores in the form of heatmap in Figure 2. For detailed scores with standard deviation, please check Table 1.
@@
From Figure 4, we can see that some models including Flux 1.1 pro and Gemini 2.0 tend not to follow the instructions, especially when the number of characters in the given prompt becomes longer. We will discuss more in Section 5.
@@ 5 Discussion
Most diffusion models demonstrate a marked decline in performance as the input text length exceeds approximately 200 characters. This threshold aligns with the 77-token limit of the CLIP text encoder, which is used in HiDream, stable diffusion series and other diffusion models. Taking a step back, as mentioned in Zhang et al. (2024a), the real effective token length of CLIP is 20 rather than 77. Considering the above two points together, it is likely that the model’s ability to capture and condition on longer instructions is restricted. Moreover, existing training corpora rarely include such long and instruction-heavy prompts, further exacerbating generalization challenges in these regimes.
@@
With longer text prompts, diffusion models increasingly fail to adhere to instruction semantics. Instead of generating a document-like image containing the target text, many models instead synthesize a naturalistic image related to the subject matter of the prompt, completely omitting any embedded text. For example, in Figure 5, when instructed to render a document stating “Asia ( , ) is the largest continent in the world by both land area and population….", some models, especially Flux 1.1 pro, return an illustration of an Asian map instead of the text itself as the number of characters increase. We hypothesize that this failure arises from three main causes: (1) the dilution of the instruction signal in long prompts, which reduces the model’s focus on the “document should contain the text:" instruction; (2) the limited capacity of CLIP-based or similar text encoders to capture long-range logical structure in extended inputs. These encoders may process such inputs more like a bag-of-words and thus compromise instruction fidelity; (3) the use of certain models, such as HiDream-L1, which incorporate LLMs (e.g., Llama-3.1-8B Grattafiori et al. (2024)) as encoders, although LLMs are primarily designed as decoders and are not well-suited for encoding tasks. Prior work Huang et al. (2024); BehnamGhader et al. (2024) has highlighted that LLMs generally underperform in encoder roles, although some modifications can improve their effectiveness.
@@
We observe consistent differences in performance across languages. English generally yields the highest accuracy, followed by French, while Chinese exhibits the lowest performance. This is likely attributable to insufficient training data for Chinese, rather than intrinsic difficulties in rendering Chinese characters. Structurally, we do not think there is a fundamental reason showing Chinese characters should be harder to generate than Latin alphabet. The performance gap underscores a need for broader multilingual data inclusion in models’ pertaining.
@@ Appendix A Appendix: Prompt Variants
Appendix B Appendix: Detailed NED Scores in table
We present the detailed NED scores and corresponding standard deviation in table 1.
Anytext 2 | TextDiffuser 2 | Qwen-Image | nano-banana | ||||
| EN | ||||||||||||||||
| 5 | ||||||||||||||||
| 10 | ||||||||||||||||
| 15 | ||||||||||||||||
| 50 | ||||||||||||||||
| 100 | ||||||||||||||||
| 200 | ||||||||||||||||
| 300 | ||||||||||||||||
| 400 | ||||||||||||||||
| 600 | ||||||||||||||||
| 800 | ||||||||||||||||
| 1000 | ||||||||||||||||
| 1500 | ||||||||||||||||
| 2000 | ||||||||||||||||
| 3000 | ||||||||||||||||
| 4000 | ||||||||||||||||
| 5000 | ||||||||||||||||
| FR | ||||||||||||||||
| 5 | ||||||||||||||||
| 10 | ||||||||||||||||
| 15 | ||||||||||||||||
| 50 | ||||||||||||||||
| 100 | ||||||||||||||||
| 200 | ||||||||||||||||
| 300 | ||||||||||||||||
| 400 | ||||||||||||||||
| 600 | ||||||||||||||||
| 800 | ||||||||||||||||
| 1000 | ||||||||||||||||
| 1500 | ||||||||||||||||
| 2000 | ||||||||||||||||
| 3000 | ||||||||||||||||
| 4000 | ||||||||||||||||
| 5000 | ||||||||||||||||
| ZH | ||||||||||||||||
| 5 | ||||||||||||||||
| 10 | ||||||||||||||||
| 15 | ||||||||||||||||
| 50 | ||||||||||||||||
| 100 | ||||||||||||||||
| 200 | ||||||||||||||||
| 300 | ||||||||||||||||
| 400 | ||||||||||||||||
| 600 | ||||||||||||||||
| 800 | ||||||||||||||||
| 1000 | ||||||||||||||||
| 1500 | ||||||||||||||||
| 2000 | ||||||||||||||||
| 3000 | ||||||||||||||||
| 4000 | ||||||||||||||||
| 5000 | ||||||||||||||||
Appendix C Appendix: CER and WER Figure
Appendix D Use of AI Tools in Manuscript Preparation
In the preparation of this manuscript, we utilized a large language model as an assistive tool. The LLM’s role was confined to improving the grammatical structure and clarity of our writing. Furthermore, it was used to assist in debugging code snippets and generating routine documentation such as docstrings. The core research ideas, experimental design, analysis, and conclusions were conceived and executed entirely by the authors. All LLM-generated outputs were critically reviewed and edited by the authors, who take full responsibility for the final content of this paper.